# Spin-glass dynamics: experiment, theory and simulation

### v1

@@ Abstract
The study of spin-glass dynamics, long considered the paradigmatic complex system, has reached important milestones. The availability of high-quality single crystals has allowed the experimental measurement of spin-glass coherence lengths of almost macroscopic dimensions, while the advent of special-purpose massive computers (the Janus Collaboration) enables dynamical simulations that approach experimental time and length scales. This review provides an account of the quantitative convergence of these two avenues of research, with precise experimental measurements of the expected scaling laws and numerical reproduction of classic experimental results, such as memory and rejuvenation. The article opens with a brief review of the defining spin-glass properties, randomness and frustration, and their experimental consequences. These apparently simple characteristics are shown to generate rich and complex physics. Models are introduced that enable quantitative dynamical descriptions, either analytically or through simulations. The many theoretical pictures of the low-temperature phase are reviewed. After a summary of the main numerical results in equilibrium, paying particular attention to the concept of temperature chaos, this review examines off-equilibrium dynamics in the absence of a magnetic field and shows how it can be related to the structure of the equilibrium spin-glass phase through the fluctuation-dissipation relations. The nonlinear response at a given temperature is then developed, including experiments and scaling in the vicinity of the spin-glass transition temperature $T_{\mathrm{g}}$. The consequences of temperature change —including temperature chaos, rejuvenation, and memory— are reviewed. The interpretation of these phenomena requires the identification of several length scales relevant to dynamics, which, in turn, generate new insights. Finally, issues for future investigations are introduced, including what is to be “nailed down” theoretically, why the Ising Edwards-Anderson model is so successful at modeling spin-glass dynamics, and experiments yet to be undertaken.
@@ I Introduction (and a little history)
The study and broad applications of spin-glass dynamics have advanced remarkably since the last major review [63], nearly forty years ago. This work aims to bring the experimental and theoretical accomplishments up to date and point to future opportunities. While many individual summaries are available [218, 296, 78, 83, 9, 281], this review combines experimental and theoretical perspectives in a synergistic format. The give and take between experiment, theory, and simulations caused by these interactions is a perfect example of the “three legs of scientific research.” The accomplishments of the Uppsala and Saclay groups were responsible for the surge of interest in the field in the late 20th and early 21st century. The following advent of powerful computational facilities custom designed for Ising spin glasses produced simulations, both in and out of equilibrium, that have enabled direct comparisons with experimental results. These in turn have inspired new experiments that have, in combination with simulations, led to a deep quantitative understanding of spin-glass dynamics.
@@ I.1 What is a spin glass?
Over the years since the original experiments of Cannella and Mydosh [74], the combination of randomness mixed with frustration [274] has been identified as the essential property of a spin glass [15]. Simply defined, frustration means that not all exchange-coupled spin orientations can be energetically satisfied for any configuration. While each property by itself results in a different class (e.g., an antiferromagnet in a triangular lattice is fully frustrated), the combination of the two is defining for spin glasses so far as we know. Spin-glass behavior is found in both metals and insulators with competing ferro- and anti-ferromagnetic couplings. The RKKY oscillating coupling [255, 165, 295] in dilute magnetic alloys [e.g., CuMn, 132, 164] and competing exchange couplings in insulators [e.g., CdCr1.7In0.3S4, 281] are stereotypical examples.
@@ I.2 Experimental milestones
Interest in the magnetic alloys now known as spin glasses began with the work of Cannella and Mydosh, which showed a cusp in the ac susceptibility of dilute Au:Fe alloys [74, 75, see Fig. 1]. The sharpness of the cusp indicated what might be a phase transition, while earlier dc magnetization measurements on similar alloys only showed a rounded transition [see, for example, 195].
Figure 1: Susceptibility for AuFe $1\leq{\text{c}}\leq 8$ at.in the region of low magnetic fields. The data were taken every $0.25$ K around the peak and every $0.5$ or 1 K elsewhere. The scatter of the points is of the order of the thickness of the lines. The open circles indicate isolated points taken at higher temperatures. Reproduced from Fig. 9 of Cannella and Mydosh [74].
A simple picture described the system as merely a distribution of independent magnetic domains, so that the cusp was an artifact, caused by the freezing of a portion of the sample at a given temperature and frequency [271, 272, 119]. In such a case, a variation in the measurement frequency would significantly change the temperature of the cusp. This question was addressed by the UCLA group [92, 91]. They measured the frequency dependence of the cusp temperature and found the maximum shift to be of the order of 50 mK from 16 Hz to 2.8 MHz (Fig. 2). This was strong evidence that the cusp was the signature of some type of phase transition.
@@
In the early years, prior to this study in an extended frequency range, the dc magnetic measurements had not revealed critical behavior. The search was on for another signature, and it seemed obvious that heat capacity should reveal some sort of thermodynamic signature. However, heat-capacity measurements by Wenger and Keesom [289, 290] showed no evidence (see Fig. 3).
Figure 3: Specific heat of Au0.92Fe0.08 in the temperature region 3-50 K is shown as a plot of $C/T$ vs $T^{2}$. The solid curve is the calculated non-magnetic contribution to the specific heat of the alloy between 0 and 30 K. The inset shows the susceptibility results of the same sample, which were provided by S. A. Werner. Reproduced from Fig. 2 of Wenger and Keesom [289] and Wenger and Keesom [290].
Other measurements were made in hopes of observing something indicating critical behavior. They included neutron scattering, resistivity, Hall, thermopower, ultrasound, muon spin resonance, ultrasound, and Mossbauer measurements, but none provided the same strong signature as ac susceptibility [see 119, for an in depth discussion of each of these probes]. In 1991, it became clear that a thermodynamic phase transition took place at $T_{\mathrm{g}}$ [141]. This work was followed by simulations [233, 38] that supported the conclusion that spin glasses experience a continuous phase transition.
The spin-glass transition temperature has been defined as the point at which there is an onset of irreversibility, which is seen by comparing two measurements of the spin-glass magnetization: lowering the temperature in the presence of a magnetic field [field-cooled magnetization (FC)]; or lowering the temperature in the absence of a magnetic field and then turning it on [zero-field-cooled magnetization (ZFC)]. Fig. 4 displays the difference, the so-called thermoremanent magnetization (TRM), between both protocols [102]. Experimentally, the FC magnetization has a weak time dependence, while ZFC and TRM measurements are strongly time dependent.
@@
where $t_{\mathrm{w}}$ is the waiting or aging time (see glossary in Sec. I.3). However, Paga et al. [231] showed that Eq. (1) is valid only for very small $H$ (see Sec. V.2).
@@
Figure 6: Field dependence of the thermoremanent magnetization (TRM) of AuFe 0.5 at.obtained after cooling from $T>T_{\mathrm{g}}$ to $T=1.2$ K in a magnetic field $H$ which is them removed and of the isothermal remanent magnetization (IRM) obtained when cooling in zero magnetic field, turning on the magnetic field $H$ at 1.2 K and then removing it. Reproduced from Fig. 4 of Tholence and Tournier [271].
Early on, the TRM and the isothermal remnant magnetization (IRM) were most commonly measured [for example, 271, see Fig. 6]. An IRM measurement consists of cooling to a temperature below the glass temperature in zero applied magnetic field. A magnetic field is applied and removed with the TRM being the remaining magnetization.
@@
$S(t,t_{\mathrm{w}};H)=(\pm){\frac{\text{d}M(t_{\mathrm{w}},T;H)}{\text{d}\ln t}}\,,$ (2)
where $+$ (--) refers to ZFC (TRM) measurements. This was first developed by Nordblad et al. [228], who showed that the time for $S(t,t_{\mathrm{w}};H)$ to peak was a useful measure of a spin-glass response time, $t_{\mathrm{w}}^{\text{eff}}$ (see Fig. 7). Typically, $t_{\mathrm{w}}^{\text{eff}}\approx t_{\mathrm{w}}$, but there are important exceptions [174] for temperatures close to $T_{\mathrm{g}}$.
Figure 7: Relaxation rates $S(t)$ of the remanent magnetization of CuMn 5 at.($T_{\mathrm{g}}=28$ K) at $T=21$ K for different waiting times $t_{\mathrm{w}}$. A relaxation rate of 0.1% of the field-cooled susceptibility value is indicated on the figure. Reproduced from Fig. 2 (c) of Nordblad et al. [228].
@@
Figure 9: A plot of $N_{\text{s}}$, the number of spins participating in barrier quenching (and barrier hopping), for CuMn 6 at.vs $t_{\mathrm{w}}$ on a log scale at fixed $T=0.89\,T_{\mathrm{g}}=28$ K. The solid curve drawn through the points is the prediction for power-law dynamics [182], while the dashed curve is the prediction for activated dynamics [115, 114] with the exchange factor set equal to $T_{\mathrm{g}}$ (i.e., independent of $T$ and $t$). The two fits are equally good. Reproduced from Fig. 3 of Joh et al. [159].
Following this work, Joh et al. [159] used the field dependence of the $S(t,t_{\mathrm{w}};H)$ peaks as a function of waiting times to quantify a correlation length of the spin-glass state, $\xi(t_{\mathrm{w}},T)$. They were able to extract an explicit value for $\xi(t_{\mathrm{w}},T)$ —the $\xi_{0}$ of Fig. 8— and its growth rate as a function of $t$, $t_{\mathrm{w}}$ and $T$. Setting the number of correlated spins $N_{\text{s}}\approx\xi(t_{\mathrm{w}},T)^{3}$, they measured $\xi(t_{\mathrm{w}},T)$ as a function of $t_{\mathrm{w}}$. Their results for CuMn 6 at.are exhibited in Fig. 9. Note the different dependence upon $t_{\mathrm{w}}$ between $\xi_{0}$ and $\xi(t_{\mathrm{w}},T)^{3}$ in Figs. 8 and 9, respectively.
@@
Figure 10: Barrier distributions, $D(\Delta)$, at $T=0$, based on measurements of $S_{R}(f,T)$ for five film thicknesses: (a) 10 nm, (b) 18 nm, (c) 25 nm, (d) 40 nm, and (e) 80 nm. The bold portions fall within our measurement bandwidth. The dashed portions are an extrapolation to guide the eye. Reproduced from Fig. 3 of Harrison et al. [144].

### v2

@@ Abstract
The study of spin-glass dynamics, long considered the paradigmatic complex system, has reached important milestones. The availability of high-quality single crystals has allowed the experimental measurement of spin-glass coherence lengths of almost macroscopic dimensions, while the advent of special-purpose massive computers —the Janus Collaboration— enables dynamical simulations that approach experimental time and length scales. This review provides an account of the quantitative convergence of these two avenues of research, with precise experimental measurements of the expected scaling laws and numerical reproduction of classic experimental results, such as memory and rejuvenation. The article opens with a brief review of the defining spin-glass properties, randomness and frustration, and their experimental consequences. These apparently simple characteristics are shown to generate rich and complex physics. Models are introduced that enable quantitative dynamical descriptions, either analytically or through simulations. The many theoretical pictures of the low-temperature phase are reviewed. After a summary of the main numerical results in equilibrium, paying particular attention to the concept of temperature chaos, this review examines off-equilibrium dynamics in the absence of a magnetic field and shows how it can be related to the structure of the equilibrium spin-glass phase through the fluctuation-dissipation relations. The nonlinear response at a given temperature is then developed, including experiments and scaling in the vicinity of the spin-glass transition temperature $T_{\mathrm{g}}$. The consequences of temperature change —including temperature chaos, rejuvenation, and memory— are reviewed. The interpretation of these phenomena requires identifying several length scales relevant to dynamics, which, in turn, generate new insights. Finally, issues for future investigations are introduced, including what is to be “nailed down” theoretically, why the Ising Edwards-Anderson model is so successful at modeling spin-glass dynamics, and experiments yet to be undertaken.
@@ I Introduction (and a little history)
The study and broad applications of spin-glass dynamics have advanced remarkably since the last major review [66], nearly forty years ago. Subsequent advances have been summarized by Fischer and Hertz [132], Young [371], Castellani and Cavagna [86], Kawamura and Taniguchi [203], Mydosh [270], Charbonneau et al. [95], Altieri and Baity-Jesi [11], Vincent [352]. This work aims to bring the experiments and theoretical accomplishments of the last decade up to date and to point to future opportunities. The give and take between experiment, theory, and simulations in the field constitutes a perfect example of the “three legs of scientific research.” The accomplishments of the Uppsala and Saclay groups were responsible for the surge of interest in the field in the late 20th and early 21st century. The following advent of powerful computational facilities custom-designed for Ising spin glasses produced simulations, both in and out of equilibrium, that have enabled direct comparisons with experimental results. These in turn have inspired new experiments that have, in combination with simulations, led to a deep quantitative understanding of spin-glass dynamics.
An important theme in our discussion will be the distinction between equilibrium and nonequilibrium results. This distinction looks somewhat artificial from an experimental point of view, because a temperature exists —the so-called glass temperature $T_{\mathrm{g}}$— such that thermal equilibrium is essentially unreachable for any $T<T_{\mathrm{g}}$. As we shall discuss at length, and at variance with the open debate for other glass formers [see, e.g., 89, 67], in the case of spin glasses the physical origin of the slow dynamics below $T_{\mathrm{g}}$ is known to be a thermodynamic phase transition. It is, therefore, possible to formulate a theory for a proper spin-glass phase for $T<T_{\mathrm{g}}$ [298]. This equilibrium phase is reachable in equilibrium simulations only for samples of fairly small linear size $L$ and not at all in experiments to date. Nevertheless, the knowledge of equilibrium spin-glass structures will be important for our endeavor, because the off-equilibrium dynamics can be viewed as an endless journey toward them.
We shall compare in some detail the results obtained experimentally with the outcome of simulations for the extremely simplified Edwards-Anderson model. The conceptual framework justifying this leap will be the renormalization group and the resulting property of universality [366, 377, 296]. Given a small number of constraints (essentially the space dimension and some crucial symmetries), and provided that the correct scaling variables are employed, quantitative agreement can be reached between experimental results and the outcome of toy models representative of the appropriate universality class. In this context, the glassy coherence length, $\xi$, will be crucial to define the natural units for the parameters.
@@ I.1 What is a spin glass?
Over the years since the original experiments of Cannella and Mydosh [82], the combination of randomness mixed with frustration [344] has been identified as the essential property of a spin glass [17]. Simply defined, frustration means that not all exchange-coupled spin orientations can be energetically satisfied for any configuration. While each property by itself results in a different class (e.g., an antiferromagnet in a triangular lattice is fully frustrated), the combination of the two is defining for spin glasses so far as we know. Spin-glass behavior is found in both metals and insulators with competing ferro- and anti-ferromagnetic couplings. The RKKY oscillating coupling [320, 194, 370] in dilute magnetic alloys [e.g., CuMn, 156, 193] and competing exchange couplings in insulators [e.g., CdCr1.7In0.3S4, 352] are stereotypical examples.
@@
It is useful to quote some of the fields where these techniques are already of paramount importance. The interested reader can easily check that the list was already long 40 years ago [257], and today it is rather impressive. We will not quote all individual contributions, which would lead us to repeat here the index of Charbonneau et al. [95], without great gain. Optimization theory probably needs to be mentioned first. The importance of the spin-glass approach in this field has become paramount [for a very remarkable analysis see, for example, 151]. The theories of communication and information are very close by. Glasses (even quantum) and the theory of jamming are also important in this list. The study of polymers has seen spin-glass theory playing a crucial role. Random lasers are a relatively new subject that was very prominent in the motivation of the Nobel prize to Giorgio Parisi. Crucial applications are also found in the biological world, on many spatial and time scales: neural networks, molecules, active matter and ecological systems are among these. Lastly we quote economics and finance, where in many contexts spin-glass analogies are crucial.
@@ I.2 Experimental milestones
Interest in the magnetic alloys now known as spin glasses began with the work of Cannella and Mydosh, which showed a cusp in the ac susceptibility of dilute Au:Fe alloys [82, 83, see Fig. 1]. Others followed with similar findings [267, 268, 179, 272]. The sharpness of the cusp indicated what might be a phase transition, while earlier dc magnetization measurements on similar alloys only showed a rounded transition [see, for example, 234].
Figure 1: Susceptibility for AuFe $1\leq{\text{c}}\leq 8$ at.% in the region of low magnetic fields. The data were taken every $0.25$ K around the peak and every $0.5$ or 1 K elsewhere. The scatter of the points is of the order of the thickness of the lines. The open circles indicate isolated points taken at higher temperatures. Reproduced from Fig. 9 of Cannella and Mydosh [82].
A simple picture described the system as merely a distribution of independent magnetic domains, so that the cusp was an artifact, caused by the freezing of a portion of the sample at a given temperature and frequency [341, 342, 140]. In such a case, a variation in the measurement frequency would significantly change the temperature of the cusp. This question was addressed by the UCLA group [107, 106]. They measured the frequency dependence of the cusp temperature and found the maximum shift to be of the order of 50 mK from 16 Hz to 2.8 MHz (Fig. 2). This was strong evidence that the cusp was the signature of some type of phase transition. In their work, a small frequency dependence, of the order of 50 mK over three decades in frequency, was observed in the most extensively studied 3000 ppm AgMn sample. At the time it was believed to be associated with critical slowing down [332] as was observed in many second-order magnetic transitions. Further work by Tholence [340] found in more concentrated samples a more significant frequency dependence (see Fig. 3). This was also the first use of $\Delta T_{\text{g}}/\Delta\ln\,\nu$ as a measure of the shift of $T_{\text{g}}$ with frequency. This frequency dependence is yet another defining characteristic of a spin-glass transition (see below the first of the Mydosh criteria for distinguishing a spin glass).
@@
Figure 3: The absolute variation of $T_{\text{f}}$ (equivalent to what we call $T_{\mathrm{g}}$ in this review) per decade of time, reported as a function of $x$, the percentage fraction of Mn, for CuMn ($1\%<x<10\%$). It can be represented by an $x^{2/3}$ law. Reproduced from Fig. 2 of Tholence [340].
In the early years, prior to this study in an extended frequency range, the dc magnetic measurements had not revealed critical behavior. The search was on for another signature. It seemed obvious that heat capacity should reveal some sort of thermodynamic signature. However, heat-capacity measurements by Wenger and Keesom [364, 365] showed no evidence (see Fig. 4).
Figure 4: Specific heat of Au0.92Fe0.08 in the temperature region 3-50 K is shown as a plot of $C/T$ vs $T^{2}$. The solid curve is the calculated nonmagnetic contribution to the specific heat of the alloy between 0 and 30 K. The inset shows the susceptibility results of the same sample, which were provided by S. A. Werner. Reproduced from Fig. 2 of Wenger and Keesom [364] and Wenger and Keesom [365].
Other measurements were made in hopes of observing something indicating critical behavior. They included neutron scattering, resistivity, Hall, thermopower, ultrasound, muon spin resonance, ultrasound, and Mossbauer measurements, but none provided the same strong signature as the ac susceptibility [see 140, for an in depth discussion of each of these probes]. In 1991, it became clear that a thermodynamic phase transition had taken place at $T_{\mathrm{g}}$ [165]. This work was followed by simulations [290, 40] that supported the conclusion that spin glasses experience a continuous phase transition.
A recent compendium of new spin-glass materials can be found in the review paper of Mydosh [270]. He introduced four criteria to distinguish a “proper” spin glass. Quoting in part here:
  1. 1.
    A tiny shift $\Delta T_{\text{g}}/(T_{\text{g}}\ln\omega)$ means a strongly interacting cooperative freezing spin glass.
  2. 2.
    For the canonical spin glass, the irreversible temperature, where the first deviation between ZFC and FC occurs, coincides with the $T_{\text{g}}$ from the cusp (see below for the definition of these two protocols).
  3. 3.
    The magnetic specific heat breaks away from its linear-in-$T$ dependence at $T_{\text{g}}$ followed by a broad maximum 20%–40% above $T_{\text{g}}$.
  4. 4.
    The waiting-time effect [see Sec. IV] is a unique characteristic of the spin-glass phase.
Unfortunately, the last part of criterion 2 is seldom the case. The cusp is most often rounded, so that the spin-glass transition temperature is usually defined by the first part of 2, namely the point at which there is an onset of irreversibility seen by comparing two measurements of the spin-glass magnetization: lowering the temperature in the presence of a magnetic field [field-cooled magnetization (FC)]; and lowering the temperature in the absence of a magnetic field and then turning it on [zero-field-cooled magnetization (ZFC)]. Fig. 5 displays the difference, the so-called thermoremanent magnetization (TRM), between both protocols [117]. Experimentally, the FC magnetization has a weak time dependence, while ZFC and TRM measurements are strongly time dependent. This irreversibility, the difference between the FC and ZFC magnetization, was first reported by Ferré et al. [130], Chamberlin et al. [93], Monod et al. [259] and Tholence and Tournier [341].
@@
where $t_{\mathrm{w}}$ is the waiting or aging time (see glossary in Sec. I.3). However, Paga et al. [288] showed that Eq. (1) is valid only as $H\rightarrow 0$ (see Sec. V.2).
@@
Figure 7: Field dependence of the thermoremanent magnetization (TRM) of AuFe 0.5 at.% obtained after cooling from $T>T_{\mathrm{g}}$ to $T=1.2$ K in a magnetic field $H$ which is them removed and of the isothermal remanent magnetization (IRM) obtained when cooling in zero magnetic field, turning on the magnetic field $H$ at 1.2 K and then removing it. Reproduced from Fig. 4 of Tholence and Tournier [341].
Early on, the TRM and the isothermal remnant magnetization (IRM) were most commonly measured [for example, 341, see Fig. 7]. An IRM measurement consists of cooling to a temperature below the glass temperature in zero applied magnetic field. A magnetic field is applied and removed with the TRM being the remaining magnetization.
@@
$S(t,t_{\mathrm{w}};H)=(\pm){\frac{\text{d}M(t,t_{\mathrm{w}};H)}{\text{d}\ln t}}$}\,, (2)
where ++ (-) refers to ZFC (TRM) measurements. This was first developed by Nordblad et al. [283], who showed that the time for $S(t,t_{\mathrm{w}};H)$ to peak was a useful measure of a spin-glass response time, $t_{\mathrm{w}}^{\text{eff}}$ (see Fig. 8). Typically, $t_{\mathrm{w}}^{\text{eff}}\approx t_{\mathrm{w}}$, but there are important exceptions [208] at temperatures close to $T_{\mathrm{g}}$.
Figure 8: Relaxation rates $S(t)$ of the remanent magnetization of CuMn 5 at.% ($T_{\mathrm{g}}=28$ K) at $T=21$ K for different waiting times $t_{\mathrm{w}}$. A relaxation rate of 0.1% of the field-cooled susceptibility value is indicated on the figure. Reproduced from Fig. 2 (c) of Nordblad et al. [283].
@@
Figure 10: A plot of $N_{\text{s}}$, the number of spins participating in barrier quenching (and barrier hopping), for CuMn 6 at.% vs $t_{\mathrm{w}}$ on a log scale at fixed $T=0.89\,T_{\mathrm{g}}=28$ K. The solid curve drawn through the points is the prediction for power-law dynamics [216], while the dashed curve is the prediction for activated dynamics [136, 135] with the exchange factor set equal to $T_{\mathrm{g}}$ (i.e., independent of $T$ and $t$). The two fits are equally good. Reproduced from Fig. 3 of Joh et al. [188].
Following this work, Vincent et al. [353] proposed a scaling law for the extraction of a length scale from the field dependence of the $S(t,t_{\mathrm{w}};H)$ peak as a function of the waiting time $t_{\mathrm{w}}$. Joh et al. [188] made use of this scaling law to extract explicit values of the correlation length of the spin-glass state $\xi(t_{\mathrm{w}},T)$ —the $\xi_{0}$ of Fig. 9— and its growth rate as a function of $t$, $t_{\mathrm{w}}$, and $T$ for the metallic spin glass CuMn 6 at.% and the insulating spin glass CdCr1.7In0.3S4. These authors implicitly assumed a nonfractal geometry for the bulk of the spin-glass domains by setting the number of correlated spins $N_{\text{s}}\approx\xi(t_{\mathrm{w}},T)^{3}$ to measure $\xi(t_{\mathrm{w}},T)$ as a function of $t_{\mathrm{w}}$. Their results for CuMn 6 at.% are exhibited in Fig. 10. Note the different dependence upon $t_{\mathrm{w}}$ between $\xi_{0}$ and $\xi(t_{\mathrm{w}},T)^{3}$ in Figs. 9 and 10, respectively. We now know that the number of correlated spins is actually $N_{\text{s}}=\xi(t_{\mathrm{w}})^{D-\theta/2}$, where $\theta$ is the so-called replicon exponent —see, e.g., Marinari et al. [242, 241] for the meaning of the replicon, and Sec. II.6 for the current method of extraction of $\xi$ from experiment.
@@
Figure 11: Barrier distributions, $D(\Delta)$, at $T=0$, based on measurements of $S_{R}(f,T)$ , the spectral density of the resistance fluctuations as a function of frequency $\it f$ and temperature $T$, for five film thicknesses: (a) 10 nm, (b) 18 nm, (c) 25 nm, (d) 40 nm, and (e) 80 nm. The bold portions fall within our measurement bandwidth. The dashed portions are an extrapolation to guide the eye. Reproduced from Fig. 3 of Harrison et al. [168].